9. Given the following data, determine the nu- merical value of each of the expressions below. Person (i) X, 610 99/ 4 5 25 4 0 7 49 749 8 i.ΣΧ Xi 10. Given the data from Exercise 9 for X and Yan that c, a constant, equals 3, determine the nu merical value for the following expressions:
Answer #1

Question 9 :

Person Xi Yi Xi +Yi XiYi Xi2 Yi2
1 6 9 15 54 36 81
2 4 5 9 20 16 25
3 6 3 9 18 36 9
4 0 7 7 0 0 49
5 7 8 15 56 49 64
Total 23 32 55 148 137 228

g) Σχβ = 137

h) (ΣΧ) =232-529

i) sum X= sum_{i=1}^{5}Xi = 23

j)  5 yī2-228

k) Σή, ΣΥ).322-1024 y = 322-1024

l) sum X +sum Y= 23 +32 = 55

Question 10 :

c= 3

Person Xi Yi C(Xi+Yi) Cxi Cyi C*Xi * Yi
1 6 9 45 18 27 162
2 4 5 27 12 15 60
3 6 3 27 18 9 54
4 0 7 21 0 21 0
5 7 8 45 21 24 168
Total 23 32 165 69 96 444

a) sum_{i=1}^{5}cXi = 69

b) sum cXY= sum_{i=1}^{5}cXiYi =444

c)sum c(X+Y)=sum_{i=1}^{5}c(Xi+Yi) =165

d) sum_{i=1}^{5}c=sum_{i=1}^{5}3= 3 *5 = 15

e) sum_{i=1}^{3}(X+c) = sum_{i=1}^{3}Xi + sum_{i=1}^{3} 3= ( 6+ 4 +6 ) + ( 3*3) = 25